For a polynomial p(x), the value of p(3) is −2. Which of the following must be true about p(x)?

A) x−5 is a factor of p(x).
B) x−2 is a factor of p(x).
C) x+2 is a factor of p(x).
D) The remainder when p(x) is divided by x−3 is −2.
1
Expert's answer
2020-04-05T13:13:40-0400

Supose, we can factorize the polynomial as the following:

"p_n(x) = (x - 3)q_{n-1} + b,"

where "q_{n-1}" is the polinomial of lower order and "b" is the reminder.

Then, according to the initial condition:

"p_n(3) = 0\cdot q_{n-1} + b = -2.\\"

Thus, the reminder "b = -2" when p(x) is divided by x−3.


Answer. D.